# : Towards Semantic Steganography via Large Language Models

### v1

@@ 4.3 Training of
Trainable parameters. To preserve the knowledge of the LLM, we use LoRA [hu2022lora] for fine-tuning. Additionally, because we introduce new tokens into the vocabulary (e.g. ), we also set the LLM token embeddings (), the LLM head (), and the projection layer trainable. By the way, the and are also trainable.
Stage 1: pre-training for steganography. As shown in FigureΒ 4, the training process consists of two stages. In the first stage, we train the LLM with LoRA, Token-to-Patch MLP , and Patch-to-Token MLP . In this stage, we DO NOT introduce the cover image. We wrap the secret message with and to construct the input sequence . Given the embedding prompt along with , we feed them into the LLM, which outputs a response token sequence , which can be formulated as:
@@
Stage 2: fine-tuning in stego image. In the second stage, we FREEZE the LLM and the , TRAIN the , and introduce the cover image . After obtaining , we reshape it from patch representation to the image representation , and add it to to obtain stego image . The entire process can be formulated as follow:
@@
Training objects. The is trained using the text generation loss and the embedding loss . The overall objective of each stage is defined as the weighted sum of these losses:
Specifically, is the auto-regressive cross-entropy (CE) loss for text generation, and is the L1 loss, which encourages the model to introduce minimal artifacts in the cover image. Those losses can be formulated as:
@@ 5.1 Experimental Settings
Model variants. is a universal framework compatible with any pre-trained LLM. In this paper, we instantiate it with four models: Llama3.2-1B [grattafiori2024llama], Qwen2.5-0.5B [qwen2025qwen25technicalreport], Gemma3-1B [team2025gemma], and MiniCPM-1B [hu2024minicpm].
Implementation details. We employ AdamW optimizer [loshchilov2017decoupled]. The weights of the text generation loss and the embedding loss are set to 1.0 and 1.0, respectively. More details can be found on the Appendix B.1.
Datasets. For the secret message, we construct the training dataset using the CNN News Stories dataset [cnn_stories_dataset], news headlines [see2017get, ao-etal-2021-pens], and paper titles [farhangi2022protoformer]. For the cover image, we use the COCO training dataset [lin2014microsoft]. Please refer to Appendix B.2 for detailed information about training datasets. We evaluate the performance of steganographic models on the sentence-to-image steganography task using the IVT dataset. In experiments, we use the COCO test dataset as the cover image dataset. For more details about IVT, please refer to Appendix A.
Evaluation metrics. To measure the quality of the stego image, we use PSNR and SSIM to evaluate embedding performance. To assess recovery performance, we use word error rate (WER), BLEU-4 [papineni2002bleu], and ROUGE-L [lin2004rouge], and BERT-Score [zhang2019bertscore] to evaluate the semantic similarity between the ground truth and the recovery message.
Benchmarks. We compare with different methods, including DwtDct [Chen2009DigitalWA], StegaStamp [tancik2020stegastamp] and LanNet [lan2023robust]. We train StegaStamp and LanNet from scratch.
@@ 5.2 Sentence-to-Image Steganography Results
Sentence-to-Image results are shown in Table 1. Notably, while existing methods struggle with this task, our demonstrates strong performance.
Results of existing methods. Notably, existing methods are not suitable for the sentence-to-image steganography task, which fundamentally differs from watermarking in that it requires encoding a substantial amount of semantic information. As shown in TableΒ 1, StegaStamp (512-bit capacity) fails to recover secret messages in IVT-S, while DwtDct (768-bit capacity) performs well only on IVT-S but fails to recover messages in IVT-M, which exceed its capacity, not to mention handling IVT-L. However, LanNet (1024-bit capacity) also fails on IVT-L.
@@
On the IVT-S, all four variants demonstrate competitive performance compared to traditional baselines in secret recovery while significantly outperforming them in visual quality of stego image. Specifically, -MiniCPM-1B achieves the best overall performance, while -Qwen2.5-0.5B, -Llama3.2-1B, and -Gemma3-1B also deliver competitive results.
On the IVT-M, which involves longer and more complex message, -MiniCPM-1B maintains superior performance across most evaluation metrics. -Qwen2.5-0.5B and -Llama3.2-1B show slight performance degradation, while -Gemma3-1B demonstrates inferior semantic accuracy despite achieving high image quality.
On the IVT-L, all models exhibit performance degradation compared to their results on IVT-M. -Qwen2.5-0.5B get the top performance, -MiniCPM-1B and -Llama3.2-1B demonstrate moderate performance declines, and -Gemma3-1B exhibits the lowest semantic accuracy.
Experimental results demonstrate that all variants achieve strong performance on both IVT-S and IVT-M. However, for these models, IVT-L remains challenging, which comprises multi-sentence structures characterized by complex syntactic patterns and diverse lexical content. More details can be found in Appendix E.
| IVT-S | IVT-M | IVT-L | ||||||||||||||||
| StegaStamp [tancik2020stegastamp] | 0.452 | 0.241 | 0.138 | 0.347 | 35.6 | 0.927 | - | - | - | - | - | - | - | - | - | - | - | - |
| DwtDct [chen2006dwt] | 0.007 | 0.699 | 0.996 | 0.996 | 31.8 | 0.898 | 0.701 | 0.490 | 0.713 | 0.829 | 27.7 | 0.776 | - | - | - | - | - | - |
| LanNet [lan2023robust] | 0.004 | 0.706 | 0.997 | 0.996 | 42.1 | 0.951 | 0.244 | 0.698 | 0.836 | 0.857 | 41.3 | 0.949 | 0.824 | 0.031 | 0.288 | 0.630 | 40.9 | 0.941 |
| S2LM-Qwen2.5-0.5B | 0.045 | 0.619 | 0.873 | 0.919 | 52.1 | 0.997 | 0.050 | 0.906 | 0.944 | 0.952 | 48.0 | 0.992 | 0.112 | 0.840 | 0.901 | 0.913 | 39.0 | 0.951 |
| S2LM-MiniCPM-1B | 0.046 | 0.635 | 0.923 | 0.959 | 59.8 | 0.999 | 0.061 | 0.913 | 0.962 | 0.965 | 52.6 | 0.996 | 0.197 | 0.762 | 0.859 | 0.889 | 46.9 | 0.988 |
| S2LM-Llama3.2-1B | 0.049 | 0.616 | 0.874 | 0.927 | 53.2 | 0.998 | 0.080 | 0.840 | 0.886 | 0.908 | 48.0 | 0.993 | 0.171 | 0.653 | 0.743 | 0.798 | 40.8 | 0.973 |
| S2LM-Gemma3-1B | 0.100 | 0.498 | 0.746 | 0.834 | 53.2 | 0.998 | 0.141 | 0.745 | 0.832 | 0.871 | 51.7 | 0.997 | 0.254 | 0.678 | 0.795 | 0.840 | 44.8 | 0.984 |
Qualitative results. As depicted in Figure 5, we provide a visualization of stego images and recovery results of -Qwen2.5-0.5B in different length of information. We mark the difference in red. As we can see, the recovery message and the ground truth keep high semantic similarity, while the stego image still has a good visual quality. More samples can be found in Appendix I.
| Capacity | Compress Ratio | Secret / Recovery | Cover / Stego | ||||
@@
| 32 Token | 1:2 | 0.0331 | 0.9470 | 0.9752 | 0.9711 | 34.61 | 0.8759 |
| 64 Token | 1:1 | 0.0353 | 0.9523 | 0.9779 | 0.9694 | 31.64 | 0.7735 |
| 128 Token | 2:1 | 0.0855 | 0.8920 | 0.9476 | 0.9347 | 33.57 | 0.8536 |
| 256 Token | 4:1 | 0.4508 | 0.4847 | 0.7074 | 0.7454 | 29.53 | 0.5802 |
5.3 Capacity of the Cover Image
In this section, we conduct empirical experiments to investigate the capacity of the cover image under the framework. We analyze the capacity in terms of tokens, setting the secret message length to 32, 64, 128, and 256 tokens, while fixing the cover image resolution at . We train -Qwen2.5-0.5B from scratch for each setting. The cover image is divided into 64 patches, which is equal to the number of secret message embeddings (SMEs). For clearer analysis, we define the compression ratio as the number of SMEs divided by the number of patches.
| Two-Stage Training | Prompt | Embedding Loss | Mask Strategy | Secret Text / Recovery Text | Cover / Stego | ||||
| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||||
| 0.3162 | 0.5925 | 0.6781 | 0.6816 | 34.13 | 0.8156 | ||||
| β | 0.2210 | 0.6819 | 0.7403 | 0.8051 | 32.93 | 0.8460 | |||
@@ 5.3 Capacity of the Cover Image
As illustrated in TableΒ 2, the model maintains strong decoding performance at 32 and 64, 128 tokens, with BERT-Scores above 0.9. However, at 256 tokens, frequent decoding errors emerging. These results indicate that while the model has effective information compression ability, its capacity saturates beyond a certain threshold, leading to a noticeable degradation in recovery quality.
Based on these empirical findings, we propose a practical heuristic: each image patch can stably carry approximately two tokens without significant loss in decoding fidelity. This guideline provides an interpretable measure of capacity for future designs and scaling strategies, offering a concrete reference point for balancing payload length and image quality in sentence-to-image steganography.
@@ 5.4 Ablation Study
Effectiveness of two-stage training. As shown in Table 3, without two-stage training strategy, the struggles to recover the secret message. That may be because introducing the image modality at an early stage disrupts the LLMβs ability to learn steganographic mechanisms.
Effectiveness of the prompt. In this experiment, we do not use prompts to guide the LLM during the embedding and decoding processes. During embedding, we simply tokenize the secret message and input it into the LLM to obtain SMEs. Similarly, during decoding, we directly feed the features extracted by the Patch-to-Token MLP into the LLM to recover the original message. As shown in TableΒ 3, this approach leads to performance improvements of 7.3%, 7.5%, and 3.0% in BLEU-4, ROUGE-L, and BERT-Score, respectively. These results confirm that the prompt effectively guides the LLM in understanding and executing the sentence-to-image task.
Effectiveness of embedding loss. In the TableΒ 3, we show the contribution of the embedding loss . The is used to encourage the LLM to produce minimal artifacts in the cover image and is applied only in the first stage. However, experiment result shows that the introduction of the can improve the embedding quality with improvement of 15.61 dB and 0.165 in the PSNR and SSIM.
Effectiveness of mask strategy. In decoding process of the first stage, we mask some input tokens to encourage LLM distribute the information in the whole cover image. As shown in FigureΒ 6, the SMEs are evenly distributed after training with the mask strategy, and significant improvements in recovery quality are observed in TableΒ 3.
@@ 6 Conclusion
In this work, we propose a new steganographic task, sentence-to-image steganography, which is an instance of semantic steganography. Additionally, we present an evaluation benchmark, Invisible Text (IVT), comprising a diverse set of sentence-level texts drawn from existing datasets and generated by large language models (LLMs). Finally, we present , a novel steganographic framework that leverages LLMs to embed sentence-level information into images. We hope our work sheds new light on the advancement of semantic steganography.
Appendix A Invisible Text Dataset
@@
A.1 IVT-S

### v2

@@ 4.3 Training of
Trainable Parameters. To preserve the knowledge of the LLM, we use LoRA [26] for fine-tuning. Additionally, because we introduce new tokens into the vocabulary (e.g. ), we also set the LLM token embeddings () and the LLM head () trainable. By the way, the and are also trainable.
| Methods | IVT-S | IVT-M | IVT-L | |||||||||||||||
| Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | |||||||||||||
| StegaStamp [60] | 0.452 | 0.241 | 0.138 | 0.347 | 35.6 | 0.927 | - | - | - | - | - | - | - | - | - | - | - | - |
[…]
| FPGP [78] | 0.150 | 0.882 | 0.925 | 0.914 | 43.5 | 0.954 | 0.758 | 0.462 | 0.539 | 0.538 | 35.6 | 0.874 | - | - | - | - | - | - |
[…]
Stage 1: Pre-Training for Steganography. As shown in Fig.Β 4, the training process consists of two stages. In the first stage, we train the LLM (with LoRA), together with and . In this stage, we DO NOT introduce the cover image. We wrap the secret message with and to construct the input sequence :
Given the embedding prompt along with , we feed them into the LLM, which outputs a response token sequence , which can be formulated as:
@@
Stage 2: Fine-Tuning on Stego Images. In the second stage, we FREEZE the LLM and , TRAIN , and introduce the cover image . After obtaining , we reshape it from patch representation to the image representation , and add it to to obtain stego image . The entire process can be formulated as follows:
@@
Training Objectives. The is trained using the text generation loss and the embedding loss . The overall objective of each stage is defined as the weighted sum of these losses:
Specifically, is the cross-entropy (CE) loss for text generation, and is the L1 loss, which encourages the model to introduce minimal artifacts in the cover image. Those losses can be formulated as:
@@ 5.1 Experimental Settings
Model Variants. is a universal framework compatible with any pre-trained LLM. In this paper, we instantiate it with four models: Llama3.2-1B [21], Qwen2.5-0.5B [55], Gemma3-1B [61], and MiniCPM-1B [28]. We chose edge-deployable language models with parameters fewer than 1B, which guarantees lightweight operation while maintaining the scalability to larger models.
Implementation Details. We employ AdamW optimizer [44]. The weights and are set to 1.0 and 1.0, respectively. More details can be found in Sec.Β 8.2 of the supplementary material.
Datasets. For the secret message, we use Wanjuan 1.0 [25], a comprehensive corpus designed for large language model pre-training. We randomly select 320,000 samples with fewer than 1,000 words as secret messages for training. For the cover image, we use the COCO training dataset [41]. Please refer to the supplementary material for detailed information about the training dataset. We evaluate the performance of steganographic models on the sentence-to-image steganography task, using the IVT dataset as secret messages and the COCO test dataset as cover images. For more details about IVT, please refer to Sec.Β 7 in the supplementary material.
Evaluation Metrics. The performance of the steganographic model is evaluated from three complementary perspectives: recoverability, imperceptibility, and security. For recoverability, we evaluate the recovery text using Word Error Rate (WER), BLEU-4 [47], ROUGE-L [40], and BERT-Score [77], which jointly assess lexical accuracy and semantic consistency. For imperceptibility, we adopt PSNR and SSIM to measure the visual similarity between the cover and stego images. For security, we employ state-of-the-art steganalysis tools including SRNet [8], SiaSigNet [71], and DBS2Net [27], to quantify the detectability of stego images.
@@ 5.2 Sentence-to-Image Steganography Results
Sentence-to-Image results are shown in Tab.Β 1. Notably, while existing methods struggle with this task, our demonstrates strong performance.
Results of Existing Methods. Existing methods are not suitable for the sentence-to-image steganography task, which fundamentally differs in that it requires encoding a substantial amount of semantic information. As shown in Tab.Β 1, StegaStamp fails to recover secret messages in IVT-S, while DwtDct and FPGP perform well only on IVT-S, but fail to recover messages in IVT-M, not to mention handling IVT-L. However, LanNet and CRMark also fail on IVT-L.
To ensure sufficient capacity for the IVT dataset, we retrain all of the baselines and increase their steganographic capacity for each IVT dataset. However, we observe a severe mode collapse problem, which is consistent with the phenomenon reported in [52]. We provide a detailed analysis in Sec.Β 12.2 of the supplementary material. This issue primarily stems from limitations in the architectural design and bitstream representation of secret information. As a result, these models are restricted to hiding bit-level information and cannot handle semantic-level content, such as sentences.
@@
On the IVT-S, all four variants demonstrate competitive performance compared to traditional baselines in secret recovery while significantly outperforming them in the visual quality of the stego image. Specifically, -MiniCPM-1B achieves the best overall performance, while -Qwen2.5-0.5B, -Llama3.2-1B, and -Gemma3-1B also deliver competitive results.
| Capacity (Token) | Compress Ratio | Secret / Recovery | Cover / Stego | ||||
@@
| 32 | 1:2 | 0.033 | 0.947 | 0.975 | 0.971 | 34.6 | 0.875 |
| 64 | 1:1 | 0.035 | 0.952 | 0.977 | 0.969 | 31.6 | 0.773 |
| 128 | 2:1 | 0.085 | 0.892 | 0.947 | 0.934 | 33.5 | 0.853 |
| 256 | 4:1 | 0.214 | 0.725 | 0.860 | 0.858 | 32.4 | 0.809 |
| 512 | 8:1 | 0.875 | 0.237 | 0.474 | 0.708 | 31.5 | 0.770 |
On the IVT-M, which involves longer and more complex messages, our variants significantly outperform prior methods. -Qwen2.5-0.5B achieves the best recovery performance among all variants while maintaining high visual quality. -MiniCPM-1B and -Gemma3-1B exhibit comparable performance.
On the IVT-L, all models exhibit performance degradation compared to their results on IVT-M. -Qwen2.5-0.5B get the top recovery performance, -MiniCPM-1B and -Llama3.2-1B demonstrate moderate performance declines, and -Gemma3-1B exhibits the lowest semantic accuracy.
Experimental results demonstrate that all variants achieve strong performance on both IVT-S and IVT-M. However, for these models, IVT-L remains challenging, which comprises multi-sentence structures characterized by complex syntactic patterns and diverse lexical content. Additional robustness experiments demonstrate that exhibits strong resistance to various attacks while maintaining a remarkably high capacity. More details can be found in Sec.Β 12.4 of the supplementary material.
Qualitative Results. As depicted in Fig.Β 5, we provide a visualization of stego images and recovery results of -Qwen2.5-0.5B in different lengths of information. We mark the difference in red. As we can see, the recovery message and the ground truth keep a high semantic similarity, while the stego image still has a good visual quality. More samples can be found in Sec.Β 15 of the supplementary material.
| No. | Two-Stage Training | Prompt | Embedding Loss | Mask Strategy | Secret / Recovery | Cover / Stego | ||||
| WER | BLEU-4 | ROUGE-L | BERT-S | PSNR | SSIM | |||||
| 1 | 0.3162 | 0.5925 | 0.6781 | 0.6816 | 34.13 | 0.8156 | ||||
| 2 | β | 0.2210 | 0.6819 | 0.7403 | 0.8051 | 32.93 | 0.8460 | |||
@@
5.3 Capacity Analysis
In this section, we conduct empirical experiments to investigate the capacity of the cover image under the framework. We analyze the capacity in terms of tokens, setting the secret message length to 32, 64, 128, 256, and 512 tokens, while fixing the cover image resolution at . We train -Qwen2.5-0.5B from scratch for each setting. The cover image is divided into 64 patches, which is equal to the number of SMEs. For more precise analysis, we define the compression ratio as the number of message tokens divided by the number of patches.
As illustrated in Tab.Β 2, the model maintains strong decoding performance for message lengths ranging from 32 to 256 tokens, with BERT-Scores above 0.8. However, at 512 tokens, frequent decoding errors emerge. These results indicate that while the has the information compression ability, its capacity saturates beyond a certain threshold, leading to a noticeable degradation in recovery quality.
Based on these empirical findings, we propose a practical heuristic: each image patch can stably carry approximately four tokens without significant loss in decoding fidelity. This guideline provides an interpretable measure of capacity for future designs and scaling strategies, offering a concrete reference point for balancing payload and image quality in sentence-to-image steganography.
5.4 Steganalysis
Following previous works [23, 29, 31, 83, 75], we gradually increase the number of training samples to investigate the minimum number of leaking samples required to train steganalysis models for detecting stego images. We retain the SRNet, SiaStegNet, and DBS2Net with cover/stego image pairs on the COCO dataset. As shown in Fig.Β 6, with the increasing number of training samples, the still achieves lower detection accuracy. More details about the statistical steganalysis can be found in Sec.Β 9 of the supplementary material.
@@ 5.5 Ablation Study
Effectiveness of Two-Stage Training. Without the two-stage training strategy, the struggles to recover the secret message (No.1 vs. No.2, No.5 vs. No.6). That may be because introducing the image modality at an early stage disrupts the LLM to learn steganographic mechanisms.
Effectiveness of the Prompt. In this experiment, we do not use prompts to guide the LLM during the embedding and decoding processes. During embedding, we simply tokenize the secret message and input it into the LLM to obtain SMEs. Similarly, during decoding, we directly feed the features extracted by the Patch-to-Token MLP into the LLM to recover the original message. Comparing experiments No.2 and No.3 reveals that this approach leads to performance improvements of 7.3%, 7.5%, and 3.0% in BLEU-4, ROUGE-L, and BERT-Score, respectively. These results confirm that the prompt effectively guides the LLM in understanding and executing the sentence-to-image task.
Effectiveness of Embedding Loss. A comparison between experiments No.3 and No.4 reveals the contribution of . is used to encourage the LLM to produce minimal artifacts in the cover image and is applied only in the first stage. The experimental results show that the introduction of can improve the embedding quality, with gains of 15.61 dB and 0.165 in the PSNR and SSIM.
Effectiveness of Mask Strategy. In the decoding process of the first stage, we mask some input tokens to encourage the LLM to distribute information across the entire cover image. As shown by the comparison between experiments No.4 and No.6, this strategy leads to significant improvements in recovery quality. We visualize the SMEs in Sec.Β 11.1 of the supplementary material, which shows that the SMEs are evenly distributed with the mask strategy.
@@ 6 Conclusion
In this work, we propose a novel concept of semantic steganography, which aims to hide high-level semantic information in carriers and redefines the traditional paradigm of steganography. Building on this insight, we present sentence-to-image steganography as an instance, which focuses on embedding sentence-level information into images. To achieve this task, we propose , a novel steganographic framework that leverages LLMs to embed sentence-level information into images, achieving a feat that previously has been out of reach. Additionally, we present an evaluation benchmark, Invisible Text (IVT), comprising a diverse set of sentence-level texts drawn from existing datasets and generated by LLMs. We hope our work sheds new light on the advancement of semantic steganography for the entire community.
Supplementary Material
Overview of the Supplementary Material
This supplementary document provides a comprehensive extension to the main paper, organized as follows:
β’
Dataset Details (Sec.Β 7): We introduce the proposed Invisible Text (IVT) benchmark, describe the construction of its three granular subsets (IVT-S/M/L), and provide detailed statistics on their text length, bit length, domains, and licenses. We also describe the construction of the LLM-generated IVTG dataset.
β’
Experimental Details (Sec.Β 8): We provide full training configurations for and baselines, including training dataset construction, model variants, optimization settings, implementation details, and retraining protocols for comparison methods.
β’
Statistical Steganalysis (Sec.Β 9): We conduct statistical steganalysis using StegExpose to evaluate the detectability of stego images and show that our method remains highly secure against classical statistical attacks.
β’
Additional Experimental Results (Sec.Β 10): We present comprehensive supplementary experiments, including using vision-language models within , constructing and benchmarking on the LLM-generated IVTG dataset, testing non-semantic messages, analyzing decoding with the original and cross-instance LLMs, applying 8-bit quantization on , and providing additional qualitative results.
β’
Ablation Study Details (Sec.Β 11): We visualize the spatial distribution of secret message embeddings under the mask strategy to analyze how it shapes SME layouts.
β’
Analysis of Existing Methods (Sec.Β 12): We analyze how existing UTF-based bitstream representations and MLP designs inherently constrain capacity and scalability of current steganographic methods.
β’
Discussion (Sec.Β 13): We discuss the motivation behind semantic steganography, as well as the broader impacts, potential risks, and mitigation strategies associated with .
β’
Background and Preliminaries (Sec.Β 14): We clarify the terminology and related notions, such as the cover image and the stego image, used throughout this work.
7 Invisible Text Dataset
@@
| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
| Average | Max | Min | Average | Max | Min | |||||
| Paper title | arxiv-10 [16] | 592.2 | 1296 | 120 | 9.7 | 20 | 2 | 4576 | 1000 | GPL-3.0 |
| News title | PENS [3] | 506.9 | 1072 | 120 | 10.3 | 20 | 2 | 5506 | 1000 | MSR License |
| Article title | examiner [34] | 423.1 | 800 | 40 | 8.6 | 19 | 1 | 4873 | 1000 | CC0 1.0 License |
| People name | name&code [63] | 113.8 | 200 | 72 | 2.1 | 4 | 2 | 857 | 1000 | - |
| Sentence | cmn-en [32] | 247.3 | 848 | 40 | 6.2 | 18 | 1 | 1946 | 1000 | CC-BY 2.0 License |
| Dialogue | SP-S1 [1] | 202.9 | 520 | 32 | 4.5 | 9 | 1 | 1924 | 1000 | - |
| Comment | sentiment140 [20] | 486.2 | 1120 | 56 | 10.7 | 20 | 1 | 4374 | 1000 | - |
| Overall | - | 367.5 | 836.6 | 68.6 | 7.4 | 15.7 | 1.4 | 19941 | 7000 | - |
| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
| Average | Max | Min | Average | Max | Min | |||||
| Paper title | arxiv-10 [16] | 1210.0 | 1904 | 696 | 22.0 | 38 | 20 | 7776 | 1000 | GPL-3.0 |
| New title | PENS [3] | 1369.4 | 1968 | 624 | 23.0 | 41 | 20 | 8003 | 1000 | MSR License |
| News highlight | CNN and Daily mail [56] | 2303.6 | 4896 | 736 | 50.3 | 100 | 20 | 13119 | 1000 | Apache 2.0 |
| Translate text | WMT-14 [7] | 1610.1 | 4776 | 672 | 33.8 | 98 | 20 | 6760 | 1000 | - |
| Joke story | stupid joke [53] | 2042.1 | 17648 | 592 | 47.6 | 100 | 20 | 10637 | 1000 | - |
| Translate text | europarl [33] | 1618.8 | 5344 | 744 | 34.0 | 99 | 20 | 6735 | 1000 | - |
| Dialogue | South Park [1] | 1367.1 | 4040 | 688 | 31.7 | 92 | 20 | 3896 | 1000 | - |
| Comment | hate text [13] | 989.9 | 1120 | 552 | 23.4 | 33 | 20 | 6564 | 1000 | MIT License |
| Moral | moral story [22] | 1369.4 | 4728 | 680 | 32.5 | 97 | 20 | 3179 | 1000 | CC BY-NC-SA 3.0 |
| Overall | - | 1542.3 | 17648 | 552 | 33.1 | 100 | 20 | 49423 | 9000 | - |
| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
| Average | Max | Min | Average | Max | Min | |||||
| Paper abstract | arxiv-10 [16] | 8589.8 | 15360 | 2424 | 157.8 | 325 | 50 | 22705 | 1000 | GPL-3.0 |
| Wiki abstract | wiki [18] | 7840.5 | 40776 | 2216 | 157.5 | 863 | 50 | 26582 | 1000 | GFDL/CC BY-SA |
| Story | stupid joke [53] | 7285.7 | 41168 | 1920 | 166.4 | 997 | 50 | 25232 | 1000 | - |
| Story | moral story [22] | 11374.1 | 42576 | 1920 | 265.0 | 956 | 50 | 20898 | 1000 | - |
| Story | Tiny Story [15] | 6957.5 | 33192 | 2600 | 166.4 | 813 | 64 | 10348 | 1000 | - |
| User review | IMDB [49] | 10457.4 | 45736 | 2048 | 231.0 | 1000 | 50 | 34259 | 1000 | IMDB License |
| User review | yelp [4] | 5996.6 | 39664 | 1968 | 138.9 | 871 | 50 | 18029 | 1000 | Yelp License |
| Composition | asap-aes [17] | 10435.9 | 39216 | 1912 | 238.5 | 897 | 50 | 17289 | 1000 | - |
| Overall | - | 8675.3 | 45736 | 1912 | 194.5 | 1000 | 50 | 118472 | 8000 | - |
7.1 IVT-S